# Rotatable and Movable Antenna-Enabled Near-Field Integrated Sensing and Communication

### v2

@@ IV-E Convergence and Complexity Analysis
Next, we analyze the computational complexity of the proposed algorithm. For each iteration, the complexity of updating $\bm{\rho}$ and $\bm{\nu}$ are $\mathcal{O}(K)$. And as calculated in the last subsection of Section III, the complexity for updating $\mathbf{W}$ and $\mathbf{R}_{0}$ via the interior method is in order of $\mathcal{O}(K^{6.5}M^{6.5}{\text{log}}\epsilon^{-1})$ with the given solution accuracy $\epsilon$. The complexity for updating $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$ via the PSO method is in order of $\mathcal{O}\left(\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M)\right)$. Therefore, the total computational complexity of Algorithm 2 is $\mathcal{O}\left(I_{\text{AO}}(2K+K^{6.5}M^{6.5}{\text{log}}\epsilon^{-1}+\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M))\right)$ , with $I_{\text{AO}}$ denoting the resultant iteration number of the AO-based algorithm.
@@ V-A Simulation Setup
We consider a global Cartesian coordinate system with the origin located at the BS. The center of TP and RP are set as $\mathcal{C}^{\text{t}}=[0,0,5+D/2]^{\mathsf{T}}$ and $\mathcal{C}^{\text{r}}=[0,0,5-D/2]^{\mathsf{T}}$, respectively. The MAs at the BS are confined to a square moving region, which is modeled as a rectangle within the planar local Cartesian coordinate system relative to the center of TP/RP, with local coordinates ranging from $[-D/2,D/2]\times[-D/2,D/2]$. The communication users are assumed to be randomly distributed in the $x-y$ plane, specifically $d_{k}\sim\mathcal{U}(15,25)$, $\theta_{k}=\frac{\pi}{2}$, and $\phi_{k}\sim\mathcal{U}(-\frac{\pi}{2},\frac{\pi}{2})$ for $k=[1,K]$. The sensing target’s position is set as $d_{0}=10$, $\theta_{0}=\frac{\pi}{3}$, and $\phi_{0}=\frac{\pi}{4}$. The simulation parameters are detailed in Table I, unless stated otherwise.
@@
$f_{\text{c}}$ Carrier frequency 24G
@@
(a) RCRB of $d_{0}$
(b) RCRB of $\theta_{0}$
(c) RCRB of $\phi_{0}$
(d) sum-rate
Figure 2: Performance versus the iteration number
@@ V-C Performance Comparison for Different Setups
(a) RCRB of $d_{0}$
(b) RCRB of $\theta_{0}$
(c) RCRB of $\phi_{0}$
Figure 3: RCRB versus $\Gamma_{\min}$
Figure 4: Sum-rate versus $C_{\max}$
(a) RCRB of $d_{0}$
(b) RCRB of $\theta_{0}$
(c) RCRB of $\phi_{0}$
Figure 5: RCRB versus $\Gamma_{\min}$ with discrete rotations
Figure 6: Sum-rate versus $C_{\max}$ with discrete rotations
To demonstrate the individual effect and synergy of RMAs’ rotations and element position movement, we introduce six different setups for the antennas at BS, namely “Transmit & receive RMA”, “Transmit RMA & receive FPA”, “Transmit only rotation & receive FPA”, “Transmit conventional MA & receive FPA”, “Transmit 1D-rotation RMA & receive FPA”, and “Transmit FPA & receive FPA”, respectively.
@@
Fig. 4 shows the communication performance comparison for different $C_{\max}$ with $\Gamma_{\min}=6$dB. As shown, the downlink sum-rate increases as the increment of $C_{\max}$ and gradually slows down, which can be regarded as a manifestation of the CRB-rate boundary [33]. In Fig. 4, the performance gain from rotations is more significant compared to the sensing-centric design because the effective aperture loss is mitigated by the rotations of TP and RP. The downlink sum-rate of the setup with transmit & receive RMA is substantially higher than transmit FPA & receive FPA. At $C_{\max}=1.4\times 10^{-6}$, the former achieves $78.7$% performance improvement over the latter. This indicates that the proposed BS design and algorithm are desirable and appealing when the ISAC network is heavily loaded.
@@ V-D Performance for Discrete Rotations
In practical deployment, continuous rotation of RMAs is difficult to realize since RMAs need to be mechanically moved by physical devices, such as a stepper motor that can only adjust the rotations of each RMA in discrete steps, which are usually specified by the motor used [14]. As such, we employ discrete rotations for the transmit RMA & receive FPA setup in Fig. 5 and Fig. 6. Specifically, we uniformly quantize the rotation angles with $2$ to $8$ bits by quantizing the initial rotation angles and the corresponding element of the particles in Algorithm 1 and Algorithm 2. From Fig. 5(a)-5(c), it is obvious that quantization reduces the sensing performance to a certain extent. As shown in Fig. 6, quantization has a more significant impact on the reduction of communication performance. As the number of quantization bits gets reduced, the downlink sum-rate gradually decreases, even falling below the setup with transmit 1D continuous rotation RMA & receive FPA. It indicates that the quantization level of RMA discrete rotation should be chosen based on the communication rate requirement and hardware conditions in practical deployment.

### v3

@@ IV-E Convergence and Complexity Analysis
Next, we analyze the computational complexity of the proposed algorithm. For each iteration, the complexity of updating $\bm{\rho}$ and $\bm{\nu}$ are $\mathcal{O}(K)$. And as calculated in the last subsection of Section III, the complexity for updating $\mathbf{W}$ and $\mathbf{R}_{0}$ via the interior method is in order of $\mathcal{O}(K^{6.5}N_{\text{t}}^{6.5}{\text{log}}\epsilon^{-1})$ with the given solution accuracy $\epsilon$. The complexity for updating $\mathbf{Q}^{\text{t}},\mathbf{Q}^{\text{r}},\bm{\varpi}^{\text{t}}$, and $\bm{\varpi}^{\text{r}}$ via the PSO method is in order of $\mathcal{O}\left(\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M)\right)$. Therefore, the total computational complexity of Algorithm 2 is $\mathcal{O}\left(I_{\text{AO}}(2K+K^{6.5}N_{\text{t}}^{6.5}{\text{log}}\epsilon^{-1}+\tau_{\max}M(2N_{\text{t}}+2N_{\text{r}}+6+\text{log}M))\right)$ , with $I_{\text{AO}}$ denoting the resultant iteration number of the AO-based algorithm.
@@ V-A Simulation Setup
We consider a global Cartesian coordinate system with the origin located at the BS. The center of TP and RP are set as $\mathcal{C}^{\text{t}}=[0,0,5+D/2]^{\mathsf{T}}$ and $\mathcal{C}^{\text{r}}=[0,0,5-D/2]^{\mathsf{T}}$, respectively. The MAs at the BS are confined to a square moving region, which is modeled as a rectangle within the planar local Cartesian coordinate system relative to the center of TP/RP, with local coordinates ranging from $[-D/2,D/2]\times[-D/2,D/2]$. The communication users are assumed to be randomly distributed in the $x-y$ plane, specifically $d_{k}\sim\mathcal{U}(15,25)$, $\theta_{k}=\frac{\pi}{2}$, and $\phi_{k}\sim\mathcal{U}(-\frac{\pi}{2},\frac{\pi}{2})$ for $k=[1,K]$. The sensing target’s position is set as $d_{0}=10$, $\theta_{0}=\frac{\pi}{3}$, and $\phi_{0}=\frac{\pi}{4}$. The simulation parameters are detailed in Table I, unless stated otherwise. It should be noted that due to the large aperture size, the target and the users are located within the Rayleigh distance.
@@
$f_{\text{c}}$ Carrier frequency 24GHz
@@
(a) RCRB of $d_{0}$.
(b) RCRB of $\theta_{0}$.
(c) RCRB of $\phi_{0}$.
(d) Sum-rate.
Figure 2: Performance versus the iteration number.
@@ V-C Performance Comparison for Different Setups
(a) RCRB of $d_{0}$.
(b) RCRB of $\theta_{0}$.
(c) RCRB of $\phi_{0}$.
Figure 3: RCRB versus $\Gamma_{\min}$.
Figure 4: Sum-rate versus $C_{\max}$.
(a) RCRB of $d_{0}$.
(b) RCRB of $\theta_{0}$.
(c) RCRB of $\phi_{0}$.
Figure 5: RCRB versus $\Gamma_{\min}$ with discrete rotations.
Figure 6: Sum-rate versus $C_{\max}$ with discrete rotations.
To demonstrate the individual effect and synergy of RMAs’ rotations and element position movement, we introduce the following six different setups for the antennas at BS:
@@
Fig. 4 shows the communication performance comparison for different $C_{\max}$ with $\Gamma_{\min}=6$dB. As shown, the downlink sum-rate increases as the increase of $C_{\max}$ and gradually slows down, which can be regarded as a manifestation of the CRB-rate Pareto boundary [37]. In Fig. 4, the performance gain from rotations is more significant compared to the sensing-centric design because the effective aperture loss is mitigated by the rotations of TP and RP. The downlink sum-rate of the setup with transmit & receive RMA is substantially higher than transmit FPA & receive FPA. At $C_{\max}=1.4\times 10^{-6}$, the former achieves $78.7$% performance improvement over the latter. This indicates that the proposed BS design and algorithm are desirable and appealing when the ISAC network is heavily loaded.
@@ V-D Performance for Discrete Rotations
In practical deployment, continuous rotation of RMAs is difficult to realize since RMAs need to be mechanically moved by physical devices, such as a stepper motor that can only adjust the rotations of each RMA in discrete steps, which are usually specified by the motor used [15]. As such, we employ discrete rotations for the transmit RMA & receive FPA setup in Fig. 5 and Fig. 6. Specifically, we uniformly discretize the rotation angles with $2$ to $8$ bits by quantizing the optimal rotation angles obtained by Algorithm 1 and Algorithm 2. From Fig. 5(a)-5(c), it is obvious that quantization reduces the sensing performance. As the number of quantization bits decreases, the system performance exhibits a gradual degradation. With increasing $\Gamma_{\min}$, the sensing performance for target’s position information consistently declines, indicated by sustained increase of RCRB. This characterizes the Pareto boundary of sensing and communication performance under discrete transmit antenna rotation angle setup. As shown in Fig. 6, quantization has a more significant impact on the reduction of communication performance. As the number of quantization bits decreases, the downlink sum-rate also gradually decreases, even falling below the setup with transmit 1D continuous rotation RMA & receive FPA. Similarly, as $C_{\max}$ increases, the growth of the downlink sum-rate gradually slows down, which reflects the CRB-rate Pareto boundary under discrete transmit antenna rotation angle setup. Fig. 5 and Fig. 6 indicate that sufficient quantization level of RMA discrete rotation should be chosen based on the performance requirement and hardware conditions in both sensing-centric and communication-centric systems.
@@ VI Conclusion
-B Proof of Proposition 1
Let $\tilde{\mathbf{R}}_{x}$, $\tilde{\bm{\Omega}}_{1},\dots,\tilde{\bm{\Omega}}_{K}$ be an arbitrary global optimum of $\mathcal{P}_{3}$. Our objective is to construct another global optimum, denoted as $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$, such that $\mathbf{R}_{x}^{\circ}$ is positive semidefinite, and ${\bm{\Omega}}_{k}^{\circ}$ is positive semidefinite and rank-one, for $k=1,\dots,K$. To achieve this, we prove the theorem by constructing $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$ from $\tilde{\mathbf{R}}_{x}$, $\tilde{\bm{\Omega}}_{1},\dots,\tilde{\bm{\Omega}}_{K}$ with $\mathbf{R}_{x}^{\circ}=\tilde{\mathbf{R}}_{x}$, $\mathbf{w}_{k}^{\circ}=\left(\mathbf{h}_{k}^{\mathsf{T}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}\right)^{-\frac{1}{2}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}$, and ${\bm{\Omega}}_{k}^{\circ}=\mathbf{w}_{k}^{\circ}(\mathbf{w}_{k}^{\circ})^{\mathsf{H}}$, for $k=1,\dots,K$.
We now prove that $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$ is also a global optimum to $\mathcal{P}_{3}$. Since the objective function is not directly determined by $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$ while (31b) and (32b) hold for $\mathbf{R}_{x}^{\circ}=\tilde{\mathbf{R}}_{x}$, we only need to validate that $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$ is also a feasible solution to $\mathcal{P}_{3}$ by proving (33) and (34) hold.
First, for any vector $\bm{\vartheta}\in\mathbb{C}^{N_{\text{t}}\times 1}$, we have
$\bm{\vartheta}^{\mathsf{H}}(\tilde{\bm{\Omega}}_{k}-{\bm{\Omega}}_{k}^{\circ})\bm{\vartheta}=\bm{\vartheta}^{\mathsf{H}}\tilde{\bm{\Omega}}_{k}\bm{\vartheta}-\left(\mathbf{h}_{k}^{\mathsf{T}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}\right)^{-1}|\bm{\vartheta}^{\mathsf{H}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}|^{2}.$ (67)
According to the Cauchy-Schwarz inequality, we can derive
$\left(\mathbf{h}_{k}^{\mathsf{T}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}\right)\left(\bm{\vartheta}^{\mathsf{H}}\tilde{\bm{\Omega}}_{k}\bm{\vartheta}\right)\geq|\bm{\vartheta}^{\mathsf{H}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}|^{2},$ (68)
so $\bm{\vartheta}^{\mathsf{H}}(\tilde{\bm{\Omega}}_{k}-{\bm{\Omega}}_{k}^{\circ})\bm{\vartheta}\geq 0$ holds for any vector $\bm{\vartheta}\in\mathbb{C}^{N_{\text{t}}\times 1}$. Therefore, $\mathbf{R}_{x}^{\circ}=\tilde{\mathbf{R}}_{x}\succeq\sum_{k=1}^{K}\tilde{\bm{\Omega}}_{k}\succeq\sum_{k=1}^{K}{\bm{\Omega}}_{k}^{\circ}$, namely constraint (33) holds for $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$. Moreover, we can derive that
$\mathbf{h}_{k}^{\mathsf{T}}\bm{\Omega}_{k}^{\circ}\mathbf{h}_{k}^{*}=\mathbf{h}_{k}^{\mathsf{T}}\mathbf{w}_{k}^{\circ}({\mathbf{w}_{k}^{\circ}})^{\mathsf{H}}\mathbf{h}_{k}^{*}=\mathbf{h}_{k}^{\mathsf{T}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}.$ (69)
Thus
$\frac{1+\Gamma_{\min}}{\Gamma_{\min}}\mathbf{h}_{k}^{\mathsf{T}}\bm{\Omega}_{k}^{\circ}\mathbf{h}_{k}^{*}=\frac{1+\Gamma_{\min}}{\Gamma_{\min}}\mathbf{h}_{k}^{\mathsf{T}}\tilde{\bm{\Omega}}_{k}\mathbf{h}_{k}^{*}$
$\geq\mathbf{h}_{k}^{\mathsf{T}}\tilde{\mathbf{R}}_{x}\mathbf{h}_{k}^{*}+\sigma_{k}^{2}=\mathbf{h}_{k}^{\mathsf{T}}\mathbf{R}_{x}^{\circ}\mathbf{h}_{k}^{*}+\sigma_{k}^{2},$ (70)
which ensures that constraint (34) holds for $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$.
With the derivation above, it is verified that $\mathbf{R}_{x}^{\circ}$, ${\bm{\Omega}}_{1}^{\circ},\dots,{\bm{\Omega}}_{K}^{\circ}$ is also a feasible solution, and furthermore, a global optimum to $\mathcal{P}_{3}$, which completes the proof.